half the perimeter of a rectangular lot, which is 50 feet longer than it is wide, is 350 feet. what are the dimensions of the lot?
Answer by chana(69)

Perimeter of the rectangle = 2L+2B
L=B+50
half the perimeter of a rectangular lot, which is 50 feet longer than it is wide, is 350 feet
So,
.5*(2* (B+50))+(2B))= 350
4B+100=700
solving for B we get, B as 150 feet
And L= 150+50
=200 feet